# L2R: Learning to Reduce Search Space for Generalizable Neural Routing Solver

### v1

@@ 4.3 Local Solution Construction
In this subsection, we develop a learning-based local construction model to select one final node from the candidate set $\mathcal{N}_{t}$ to construct the partial solution at each step. Similar to previous work, we also use the initial node $\pi_{1}$ and the latest node $\pi_{t-1}$ to represent the current partial solution. First of all, the embedding of each node can be obtained by a shared linear transformation:
@@
where $W^{(0)}$ and $b^{(0)}$ and learnable parameters. Then we can obtain the embedding of the partial graph:
@@
In our proposed local construction model, one attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Layer Normalization (Ba et al., 2016) and skip-connection (He et al., 2016). The detailed process can be found in Section B.3. After $L$ attention layers, the final embedding $\widetilde{H}^{(L)}=[\mathbf{h}_{\pi_{1}}^{(L)},\ \mathbf{h}_{\pi_{t-1}}^{(L)},\ H_{\mathcal{N}_{t}}^{(L)}]\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ contains advanced feature representations of the initial, last, and $k$ candidate nodes.
Finally, similar to previous work (Kool et al., 2019; Kwon et al., 2020), we can compute the probabilities for selecting each node in $\mathcal{N}_{t}$ via a compatibility module:
$u^{t}_{i}=\begin{cases}\xi\cdot\text{tanh}\left(\frac{\hat{\mathbf{h}}_{(C)}^{t}\left(\mathbf{h}_{i}^{(L)}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}+a_{t-1,i}\right)&\text{if}\ i\in\mathcal{N}_{t}\\ -\infty&\text{otherwise}\end{cases},$ (9)
where $\xi$ is the clipping parameter, $\hat{\mathbf{h}}_{(C)}^{t}=\mathbf{h}_{\pi_{1}}^{(L)}+\mathbf{h}_{\pi_{t-1}}^{(L)}$, and $a_{t-1,i}=f(|\mathcal{N}_{t}|,d_{t-1,i})$ represents the adaptation bias between each candidate node and the current node, following Zhou et al. (2024). The probability of selecting the $i$-th node as next visiting node can be calculated as
$p_{\bm{\theta}}(\pi_{t}=i\mid S,\pi_{1:t-1})=\frac{e^{u_{i}^{t}}}{\sum_{j=1}^{|\mathcal{N}_{t}|}e^{u_{j}^{t}}}.$ (10)
Thereby, the probability of generating a complete solution $\mathbf{\pi}$ for an instance $S$ can be calculated as
$p_{\bm{\theta}}(\mathbf{\pi}\mid S)=\prod_{t=2}^{N}p_{\bm{\theta}}(\pi_{t}\mid S,\pi_{1:t-1}).$ (11)
More details of local construction are in Appendix B.
@@ 4.4 Training
Our proposed L2R framework has a learning-based reduction model and a local construction model, which can be trained by a joint training scheme. We denote the learnable parameters as $\bm{\theta}_{R}$ and $\bm{\theta}_{L}$ for the reduction model and local construction model respectively, and hence $\bm{\theta}=\{\bm{\theta}_{R},\ \bm{\theta}_{L}\}$ contains all learnable parameters in the L2R framework.
Following Kool et al. (2019), L2R is trained by the REINFORCE (Williams, 1992) gradient estimator:
$\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}|S)=\mathbb{E}_{p_{\bm{\theta}}(\pi|S)}\left[\left(R(\pi|S)-b(S)\right)\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\pi|S)\right],$ (12)
where $R(\pi\mid S)$ represents the total reward (e.g., the negative value of tour length) of instance $S$ given a specific solution $\pi$, and $b(S)$ is the greedy rollout baseline. The detailed training process are provided in Appendix D.
@@ 5 Experiments
In this section, we comprehensively evaluate our proposed model against classical and learning-based solvers on synthetic and real-world TSP and CVRP instances. Notably, our model is trained solely on 100-node TSP instances from uniform distribution. We assess its generalization performance on: (1) scalability to problem sizes up to 1 million nodes and (2) robustness to varying node distributions.
@@ 5.1 Experimental Setup
For all problems, we generate synthetic instances following the methodology outlined in Kool et al. (2019). Specifically, we construct TSP test datasets with uniformly distributed nodes at six scales: 1K, 5K, 10K, 50K, 100K, and 1M. Following Fu et al. (2021), the TSP1K test set consists of 128 instances, while the larger TSP datasets each contain 16 instances. For the CVRP, we adhere to the capacity constraints specified in Hou et al. (2022) and generate five test datasets with scales of 1K, 5K, 7K, 10K, and 1M. Each dataset includes 100 instances, except for CVRP10K and CVRP1M, which contain 16 instances each.
The optimal solutions for the TSP and CVRP instances are obtained using LKH3 (Helsgaun, 2017) and HGS (Vidal, 2022), respectively. To evaluate cross-distribution generalization performance, we test on the TSP/CVRP5K instances from INViT (Fang et al., 2024). Additionally, we validate the real-world performance of L2R using symmetric EUC_2D instances from TSPLIB (Reinelt, 1991) and CVRPLIB Set-XXL (Arnold et al., 2019).
@@
TSP1K TSP5K TSP10K TSP50K TSP100K
Method Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time
@@
INViT-3V greedy 24.66 (6.66%) 9.0s 54.49 (6.90%) 1.2m 76.85 (7.07%) 3.7m 171.42 (7.18%) 1.3h 242.26 (7.20%) 5.0h
L2R greedy 24.14 (4.42%) 0.05s 53.33 (4.62%) 1.6s 75.16 (4.71%) 3.5s 167.65 (4.82% ) 30s 236.60 (4.70%) 90.8s
L2R PRC100 23.61 (2.14%) 2.2s 52.34 (2.68%) 24s 73.86 (2.90%) 28s 165.02 (3.18% ) 1.5m 234.10 (3.59%) 2.6m
L2R PRC500 23.51 (1.68%) 10.9s 52.19 (2.40%) 1.7m 73.61 (2.56%) 2.1m 164.37 (2.77% ) 5.3m 232.89 (3.05%) 7.1m
L2R PRC1,000 23.49 (1.60%) 22s 52.15 (2.31%) 3.1m 73.56 (2.48%) 3.8m 164.22 (2.68% ) 10.2m 232.49 (2.88%) 13m
Table 3: Comparison on CVRP instances following the setting in Hou et al. (2022).
CVRP1K CVRP5K CVRP7K CVRP10K
Method Obj. (Time) Obj. (Time) Obj. (Time) Obj. (Time)
@@
INViT-3V greedy 48.2 (11s) 146.6 (1.4m) 197.6 (2.4m) 262.1 (4.3m)
L2R greedy 46.3 (0.1s) 137.2 (0.4s) 182.0 (0.7s) 238.1 (3.9s)
L2R PRC100 44.7 (3.6s) 133.2 (11.2s) 178.4 (15.5s) 235.2 (33.9s)
L2R PRC500 44.3 (17s) 131.7 (53.3s) 176.7 (1.2m) 233.2 (2.5m)
L2R PRC1,000 44.2 (34.2s) 131.1 (1.8m) 175.8 (2.5m) 232.3 (4.9m)
Table 4: Comparison on TSP instances with one million nodes.
TSP1M CVRP1M
Method† Obj. Gap Time Obj. Gap Time
Concorde/HGS N/A N/A N/A OOM
@@
L2R greedy 747.29 4.67% 33.67 m 17215.83 0.00$\ddagger$ 41.8m
  • Except for INViT, which preserves the original neighborhood structure, other methods adopt D-SSR to align their search spaces with that of L2R.
@@
    Because classical solvers (i.e., HGS and LKH3) fail to collect feasible solutions on CVRP1M, the gap reported for CVRP1M is measured relative to the best-obtained solution.
Model & Training Setting
For all experiments, we use an embedding dimension of $128$ and a feed-forward layer dimension of $512$. To enhance geometric pattern recognition in VRPs, we integrate scale and distance information into the attention mechanism (see Appendix C for implementation details). The local construction model employs $6$ attention layers. Consistent with Kool et al. (2019), we set the clipping parameter $\xi=10$ in Equation 9. The hyperparameter $k$ is configured as $20$ for TSP and $50$ for CVRP. All experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU (24GB memory).
Our model is exclusively trained on uniform-distributed instances with $100$ nodes. We employ the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of $\eta=10^{-4}$ and a learning rate decay of $0.98$ per epoch. Training spans $100$ epochs with $2\text{,}500$ batches per epoch. Due to memory constraints, batch sizes differ across problems—$180$ for TSP and $64$ for CVRP. Identical model configurations are maintained across all experiments. Additional implementation details are provided in Appendix D (training setting) and Appendix E (model setting).
Baseline
We compare L2R with the following methods: (1) Classical Solver: Concorde (Applegate et al., 2006), LKH3 (Helsgaun, 2017), HGS (Vidal, 2022); (2) Constructive NCO: POMO (Kwon et al., 2020), Omni_VRP (Zhou et al., 2023), ELG (Gao et al., 2024), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), and INViT (Fang et al., 2024); (3) Two-Stage NCO: TAM (Hou et al., 2022), L2D (Li et al., 2021), RBG (Zong et al., 2022), H-TSP (Pan et al., 2023), and GLOP (Ye et al., 2024).
Metrics and Inference
For most NCO baseline methods, we execute the source code provided by the authors using default settings. Results marked with an asterisk (*) are directly obtained from the corresponding papers. Some methods fail to produce feasible solutions within a reasonable time limit (e.g., several days), which is denoted by ‘N/A’. The notation ‘OOM’ indicates that the memory consumption exceeds the available memory limits. For L2R, we report two types of results—those obtained by greedy trajectory (greedy) and those derived from Parallel local ReConstruction (PRC) under different numbers of iterations (Luo et al., 2024). The parallel approach demonstrates promising results by effectively trading computing time for improved solution quality. For PRC, the initial solutions are generated using the greedy trajectory. PRC100 refers to $100$ iterations, with the longest destruction length per iteration set to $1\text{,}000$ to balance speed and effectiveness. Further details about PRC are available in Luo et al. (2024).
Table 5: Comparison on cross-distribution generalization.
TSP5K, Cluster TSP5K, Explosion TSP5K, Implosion
Method Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time
LKH3 28.84 (0.00%) -- 31.98 (0.00%) -- 45.04 (0.00%) --
ELG 35.42 (22.83%) 1.6m 38.60 (20.71%) 1.6m 52.95 (17.55%) 1.6m
Omni_VRP 44.56 (54.53%) 1.1m 48.32 (51.09%) 1.1m 67.66 (50.20%) 1.1m
INViT-3V greedy 31.32 (8.60%) 1.3m 35.68 (11.59%) 1.3m 48 87 (8.49%) 1.3m
L2R greedy 30.44 (5.56%) 1.4s 34.79 (8.78%) 1.3s 47.77 (6.06%) 1.4s
L2R PRC1,000 29.66 (2.87%) 2.3m 33.13 (3.61%) 2.3m 46.26(2.71%) 2.3m
CVRP5K, Cluster CVRP5K, Explosion CVRP5K, Implosion
Method Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time
HGS 472.44 (0.00%) -- 350.82 (0.00%) -- 582.70 (0.00%) --
ELG 558.13 (18.14%) 2.4m 397.18 (13.21%) 2.2m 626.37 (7.50%) 2.2m
Omni_VRP 576.62 (22.05%) 1.3m 466.92 (33.09%) 1.4m 816.94 (40.20%) 1.3m
INViT-3V greedy 514.97 (9.00%) 1.6m 381.14 (8.64%) 1.6m 637.03 (9.32%) 1.6m
L2R greedy 486.24 (2.92%) 1.6s 366.98 (4.61%) 1.6s 606.19 (4.03%) 1.6s
L2R PRC1,000 480.12 (1.63%) 3.5m 357.61 (1.94%) 3.5m 591.61 (1.53%) 3.5m
  • All datasets and optimal solutions are obtained from INViT(Fang et al., 2024) and each contains $20$ instances. Note that for CVRP instances, the capacity is unified at $50$.
  • $\ddagger$
    The instance augmentation technique is not employed to prevent methods from exceeding memory limits.
Table 6: Comparison on large-scale TSPLIB (Reinelt, 1991) instances ($9001,000\leq N\leq 85,$ 900) and CVRPLIB Set-XXL (Arnold et al., 2019) ($0003,000\leq N\leq 30,$ 000).
Method 1K $\leq N\leq$ 5K 5K <$N\leq$ 100K All Solved#
(23 instances) (10 instances) (33 instances)
BQ bs16 10.65% 30.58%† -- 26/33
LEHD greedy 11.14% 39.34%† -- 30/33
ELG aug×$\times$8 11.34% OOM -- 23/33
INViT-3V greedy 11.49% 10.00% 11.04% 33/33
L2R greedy 9.50% 7.23% 8.81% 33/33
Method 3K <$N\leq$ 7K 7K <$N\leq$ 30K All Solved#
(4 instances) (6 instances) (10 instances)
BQ bs16 20.20% OOM -- 4/10
LEHD greedy 22.22% 32.80%† -- 6/10
ELG aug×$\times$8 16.82% † OOM -- 2/10
INViT-3V greedy 20.74% 26.64% 24.28% 10/10
L2R greedy 13.36% 11.58% 12.29% 10/10
@@
  • $\ddagger$
    The RRC and PRC techniques are not employed due to computational efficiency.
5.2 Performance Evaluation
VRPs with Uniform Distribution

### v2

@@ 3.3 Local Solution Construction
In this subsection, we develop a learning-based local solution construction model to select one final node from the candidate set $\mathcal{N}_{t}$ to construct the partial solution at construction step $t$. Similar to previous work [16, 18, 10, 9], we also use the initial node $\pi_{1}$ and the latest node $\pi_{t-1}$ to represent the current partial solution. First of all, the embedding of each node can be obtained by a shared linear transformation:
@@
where $W^{(0)}$ and $b^{(0)}$ are learnable parameters. Then we obtain the embedding of the partial graph:
@@
In our proposed local solution construction model, one attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Layer Normalization [30] and skip-connection [31]. The detailed process can be found in Section F.3. After $M$ attention layers, the final embedding $\widetilde{H}^{(M)}=[\mathbf{h}_{\pi_{1}}^{(M)},\ \mathbf{h}_{\pi_{t-1}}^{(M)},\ H_{\mathcal{N}_{t}}^{(M)}]\in\mathbb{R}^{(2+|\mathcal{N}_{t}|)\times d}$ contains advanced feature representations of the initial, last, and $k$ candidate nodes.
Finally, we compute the probabilities for selecting each node in $\mathcal{N}_{t}$ via another compatibility module different from Equation 6:
$u^{L}_{t,i}=\begin{cases}\xi\cdot\text{tanh}\left(\frac{\hat{\mathbf{h}}_{(C)}^{t}\left(\mathbf{h}_{i}^{(M)}\right)^{\mathrm{T}}}{\sqrt{d_{k}}}+a_{t-1,i}^{L}\right)&\text{if}\ i\in\mathcal{N}_{t}\\ -\infty&\text{otherwise}\end{cases},$ (10)
where $\hat{\mathbf{h}}_{(C)}^{t}=\mathbf{h}_{\pi_{1}}^{(M)}+\mathbf{h}_{\pi_{t-1}}^{(M)}$ and $a_{t-1,i}^{L}$ represents the adaptation bias between each node $i\in\mathcal{N}_{t}$ and $\pi_{t-1}$. The probabilities $\mathbf{p}=\{p_{i}|i\in\mathcal{N}_{t}\}$ of selecting the candidate node $i$ can be calculated as
$\mathbf{p}=\mathrm{softmax}(\mathbf{u}^{L}).$ (11)
More details of the local solution construction model are provided in Appendix F.
@@ 3.4 Training
Our proposed L2R framework has a learning-based reduction model and a local solution construction model, which can be trained by a joint training scheme. We denote the learnable parameters as $\bm{\theta}_{R}$ and $\bm{\theta}_{L}$ for the reduction model and local solution construction model, respectively. Hence $\bm{\theta}=\{\bm{\theta}_{R},\ \bm{\theta}_{L}\}$ contains all learnable parameters in the L2R framework.
L2R samples the node $\pi_{t}$ by probabilities $\mathbf{p}=\{p_{i}|i\in\mathcal{N}_{t}\}$ and add it to current partial solution, a complete solution $\pi=(\pi_{1},\dots,\pi_{n})$ for an instance $S$ is constructed by $\bm{\theta}_{L}$ with $n$ decoding steps. Since the reduction model is only responsible for selecting $k$ candidate nodes and does not participate in the actual node selection process, we use the total reward $\mathcal{R}(\pi\mid S,{\bm{\theta}_{L}})$ (e.g., the negative value of tour length) of instance $S$ given a specific solution $\pi$ generated by $\bm{\theta}_{L}$ for the training of both $\bm{\theta}_{R}$ and $\bm{\theta}_{L}$ simultaneously, and we define $\tau=(\tau_{1},\dots,\tau_{n})$ as the set of sampled nodes by $\bm{\theta}_{R}$. Following previous NCO works [16, 28, 32], L2R is trained by the REINFORCE [33] gradient estimator:
$\mathcal{L}_{\mathrm{Joint}}(\bm{\theta})=\mathcal{L}_{R}(\bm{\theta}_{R})+\mathcal{L}_{L}(\bm{\theta}_{L}),$ (12)
$\nabla_{\bm{\theta}_{R}}\mathcal{L}_{R}(\bm{\theta}_{R})=\mathbb{E}_{o(\tau|S,{\bm{\theta}_{R}})}\left[\left(\mathcal{R}(\pi|S,{\bm{\theta}_{L}})-b(S)\right)\nabla_{\bm{\theta}_{R}}\log o(\tau|S)\right],$ (13)
$\nabla_{\bm{\theta}_{L}}\mathcal{L}_{L}(\bm{\theta}_{L})=\mathbb{E}_{p(\pi|S,{\bm{\theta}_{L}})}\left[\left(\mathcal{R}(\pi|S,{\bm{\theta}_{L}})-b(S)\right)\nabla_{\bm{\theta}_{L}}\log p(\pi|S)\right],$ (14)
where $o(\mathbf{\tau}\mid S)=\prod_{t=2}^{n}o(\tau_{t}\mid S,\pi_{1:t-1})$, $p(\mathbf{\pi}\mid S)=\prod_{t=2}^{n}p(\pi_{t}\mid S,\pi_{1:t-1})$, $b(S)$ is the greedy rollout baseline. The detailed training process is provided in Appendix H.
@@ 4 Experiments
In this section, we comprehensively evaluate our proposed model against classical and learning-based solvers on synthetic and real-world TSP and CVRP instances [^1] We provide detailed definitions and formulations of TSP and CVRP in Appendix A.. Notably, our model is trained solely on 100-node TSP instances from uniform distribution. We assess its generalization performance on: (1) scalability to problem sizes up to 1 million nodes and (2) robustness to varying node distributions.
@@ 4.1 Experimental Setup
For all problems, we generate synthetic instances following the methodology outlined in Kool et al. [16]. Specifically, we construct TSP test datasets with uniformly distributed nodes at six scales: 1K, 5K, 10K, 50K, 100K, and 1M. Following Fu et al. [22], the TSP1K test set consists of 128 instances, while the larger TSP datasets each contain 16 instances. For the CVRP, we adhere to the capacity constraints specified in Hou et al. [34] and generate five test datasets with scales of 1K, 5K, 7K, 10K, and 1M. Each dataset includes 100 instances, except for CVRP10K and CVRP1M, which contain 16 instances.
The optimal solutions for the TSP and CVRP instances are obtained using LKH3 [3] and HGS [4], respectively. To evaluate cross-distribution generalization performance, we test on the TSP/CVRP5K instances obtained from INViT [28]. Additionally, we validate the real-world performance of L2R using symmetric EUC_2D instances from TSPLIB [29] and CVRPLIB Set-XXL [35].
Model & Training Setting
For all experiments, we use an embedding dimension of $128$ and a feed-forward layer dimension of $512$. To enhance geometric pattern recognition in VRPs, we integrate scale and distance information into the attention mechanism (see Appendix G for implementation details). The local construction model employs $6$ attention layers. Consistent with Kool et al. [16], we set the clipping parameter $\xi=10$ in Equation 6 and Equation 10. The hyperparameter $k$ is configured as $20$ for TSP and $50$ for CVRP. All experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU (24GB memory).
Our model is exclusively trained on uniformly distributed instances with $100$ nodes. We employ the Adam optimizer [36] with an initial learning rate of $\eta=10^{-4}$ and a learning rate decay of $0.98$ per epoch. Training spans $100$ epochs with $2\text{,}500$ batches per epoch. Due to memory constraints, batch sizes differ across problems—$180$ for TSP and $60$ for CVRP. The same pre-trained model is used in all experimental evaluations for each problem. Additional implementation details are provided in Appendix H (pseudocode for training) and Appendix I (training and model settings).
Baseline
We compare L2R with the following methods: (1) Classical Solver: Concorde [37], LKH3 [3], HGS [4]; (2) Constructive NCO: POMO [18], Omni_VRP [38], ELG [26], BQ [10], LEHD [9], and INViT [28]; (3) Two-Stage NCO: TAM [34] and GLOP [39].
Metrics and Inference
For most NCO baseline methods, we execute the source code provided by the authors using default settings. Results marked with an asterisk (*) are directly obtained from the corresponding papers. Some methods fail to produce feasible solutions within a reasonable time limit (e.g., several days), which is denoted by ‘N/A’. The notation ‘OOM’ indicates that the memory consumption exceeds the available memory limits. For L2R, we report two types of results, which are obtained by greedy trajectory (greedy) and those derived from Parallel local ReConstruction (PRC) under different numbers of iterations [40]. The parallel approach demonstrates promising results by effectively trading computing time for improved solution quality. For PRC, initial solutions are generated using the greedy trajectory. PRC100 refers to $100$ iterations, with the longest destruction length per iteration set to $1\text{,}000$ to balance speed and effectiveness. Further details about PRC are available in [40].
@@
Method TSP1K TSP5K TSP10K TSP50K TSP100K
Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time
@@
INViT-3V greedy 24.67 (6.71%) 0.4s 54.46 (6.84%) 12.7s 76.87 (7.09%) 34.9s 171.42 (7.18%) 4.9m 242.26 (7.20%) 18.8m
L2R greedy 24.16 (4.49%) 0.05s 53.36 (4.69%) 1.8s 75.24 (4.82%) 4.1s 167.70 (4.86% ) 35.5s 236.81 (4.79%) 1.8m
L2R PRC100 23.62 (2.18%) 2.5s 52.41 (2.82%) 20.1s 73.95 (3.03%) 26.3s 165.16 (3.27% ) 1.8m 234.36 (3.70%) 3.2m
L2R PRC500 23.54 (1.82%) 12.6s 52.25 (2.50%) 1.6m 73.73 (2.72%) 2.0m 164.61 (2.92% ) 6.6m 233.24 (3.21%) 8.8m
L2R PRC1,000 23.52 (1.72%) 24.9s 52.20 (2.40%) 3.1m 73.66 (2.62%) 3.8m 164.41 (2.80% ) 12.4m 232.77 (3.00%) 15.5m
4.2 Performance Evaluation
Cross-Size Generalization
Table 4: Comparison on CVRP instances following the setting in Hou et al. [34].
Method CVRP1K CVRP5K CVRP7K CVRP10K
Obj. (Time) Obj. (Time) Obj. (Time) Obj. (Time)
@@ 4.2 Performance Evaluation
INViT-3V greedy 48.3 (1s) 146.2 (7s) 197.3 (11s) 262.5 (1.2m)
L2R greedy 45.8 (0.1s) 136.0 (0.5s) 180.5 (0.8s) 236.7 (4.4s)
L2R PRC100 44.7 (4.0s) 132.7 (14.4s) 177.2 (19.8s) 233.2 (39.9s)
L2R PRC500 44.4 (20.0s) 131.2 (1.2m) 175.5 (1.6m) 230.9 (3.0m)
L2R PRC1,000 44.3 (40.3s) 130.6 (2.3m) 174.7 (3.2m) 230.1 (6.0m)
[…]
Table 5: Comparison of TSP and CVRP instances with one million nodes.
Method TSP1M CVRP1M
Obj. Gap Time Obj. Gap Time
Concorde N/A N/A N/A --
HGS -- OOM
@@
L2R greedy 748.18 4.79% 1.2h 17231.85 0.00% 1.4h
Cross-Distribution Generalization
We evaluate the cross-distribution performance of L2R on TSP5K/CVRP5K instances from three distinct distributions: cluster, explosion, and implosion. As shown in Table 6, L2R consistently achieves the best performance among all comparable methods across these distributions. These results further highlight the robust generalization capabilities of our proposed L2R.
Results on Benchmark Dataset
We further assess the generalization performance of L2R on instances from CVRPLIB Set-XXL [35] and TSPLIB [29]. As demonstrated in Table 7, L2R maintains its position as the best-performing model across instances of varying scales, underscoring its practical applicability in real-world scenarios. Detailed results are provided in Appendix J.
4.3 Additional Analyses
L2R vs. D-SSR
To validate the effectiveness of our proposed L2R, we conduct a comparative experiment against a dynamic distance-based SSR model, ensuring identical experimental settings. We systematically analyze the advantages of L2R, mainly including: (1) overall performance on real-world instances, (2) optimality gap and optimality ratio per instance, (3) impacts of different candidate node selection strategies, and (4) solution visualizations comparing the two approaches. As detailed in Section K.1, L2R consistently demonstrates better candidate selection on real-world problems and significantly outperforms the dynamic D-SSR baseline. These findings further underscore the superior practical applicability of L2R in real-world scenarios.
Complexity Analysis
We present the time and space complexity analyses for three different constructive models: L2R with learning-based SSR, INViT with distance-based SSR, and LEHD without SSR. The results in Section K.2 demonstrate that our local solution construction model achieves significantly lower computational complexity than both LEHD and INViT. This advantage is particularly evident in the substantially reduced runtime compared to existing representative models.
Table 6: Comparison on cross-distribution generalization.
Method TSP5K, Cluster TSP5K, Explosion TSP5K, Implosion CVRP5K, Cluster CVRP5K, Explosion CVRP5K, Implosion
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Optimal 0.00% -- 0.00% -- 0.00% -- 0.00% -- 0.00% -- 0.00% --
ELG $\ddagger$ 22.83% 1.6m 20.71% 1.6m 17.55% 1.6m 18.14% 2.4m 13.21% 2.2m 7.50% 2.2m
Omni_VRP $\ddagger$ 54.53% 1.1m 51.09% 1.1m 50.20% 1.1m 22.05% 1.3m 33.09% 1.4m 40.20% 1.3m
INViT-3V greedy 8.20% 11.3s 11.48% 11.3s 8.52% 11.3s 9.05% 29.4s 8.44% 29.3s 8.77% 23.2s
L2R greedy 6.14% 1.5s 9.54% 1.5s 6.17% 1.5s 2.65% 1.8s 4.38% 1.8s 3.15% 1.8s
L2R PRC1,000 3.16% 2.6m 3.52% 2.6m 2.87% 2.6m 0.99% 4.2m 1.14% 4.2m 0.73% 4.2m
  • All datasets are obtained from INViT[28] and each contains $20$ instances.
@@ 4.3 Additional Analyses
    The instance augmentation technique is not employed for comparable methods to prevent methods from exceeding memory limits.
Table 7: Comparison on large-scale TSPLIB [29] instances ($9005,000\leq N\leq 85,$ 900) and CVRPLIB [35] instances ($0003,000\leq N\leq 30,$ 000).
TSPLIB CVRPLIB-XXL
Method 1K $\leq N\leq$ 5K 5K <$N\leq$ 100K All Solved# 3K $\leq N\leq$ 7K 7K <$N\leq$ 30K All Solved#
(23 instances) (10 instances) (33 instances) (4 instances) (6 instances) (10 instances)
GLOP 6.16% 7.69% 6.62% 33/33 17.07% 21.32% 19.62% 10/10
TAM-LKH3* -- -- -- -- 20.44% -- -- --
BQ bs16 10.65% 30.58%† -- 26/33 20.20% OOM -- 4/10
LEHD greedy 11.14% 39.34%† -- 30/33 22.22% 32.80%† -- 6/10
LEHD RRC1000 4.00% 18.46%† -- 30/33 14.06% 21.52%† -- 6/10
POMO aug×$\times$8 62.81% OOM -- 23/33 331.24%† OOM -- 2/10
ELG aug×$\times$8 11.34% OOM -- 23/33 16.82% † OOM -- 2/10
INViT-3V greedy 11.49% 10.03% 11.05% 33/33 20.74% 26.64% 24.28% 10/10
L2R greedy 9.16% 7.36% 8.61% 33/33 12.05% 11.22% 11.55% 10/10
L2R PRC100 3.60% 4.58% 3.89% 33/33 8.62% 9.34% 9.05% 10/10
@@
4.4 Ablation Study
We conduct a detailed ablation study and analysis to demonstrate the effectiveness and robustness of L2R, mainly including: (1) L2R vs. Distance-based SSR with Larger Search Space (see Section L.1); (2) Effects of Adaptation Bias in Compatibility (see Section L.2); (3) Comparison between Different Attention Mechanisms (see Section L.3).
5 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based Learning-to-Reduce (L2R) framework for solving large-scale vehicle routing problems. Our approach adaptively selects a small set of candidate nodes at each construction step, enabling efficient search space reduction while maintaining solution quality. Extensive experiments demonstrate that L2R, trained exclusively on 100-node instances from the uniform distribution, achieves remarkable performance on TSP and CVRP instances with up to 1 million nodes from the uniform distribution and over 80K nodes from other distributions.
Limitation and Future Work
While L2R achieves impressive performance on large-scale TSP and CVRP instances, it currently relies on a simple combined loss for training. In future work, we aim to develop a more effective loss function to further improve the training efficiency of L2R. In addition, we plan to extend the L2R framework’s applicability to tackle routing problems with complex constraints without compromising solution quality.
Appendix A Definitions and Formulations of TSP and CVRP
In this section, we provide the definitions and mathematical formulations of the Traveling Salesman Problem (TSP) and the Capacitated Vehicle Routing Problem (CVRP), thereby enabling readers to obtain a clearer and more profound understanding of the fundamental nature of these problems.
A.1 Traveling Salesman Problem (TSP)
The TSP is one of the most representative combinatorial optimization problems, and we focus on the 2D Euclidean TSP in this paper [14, 16]. Given an instance $S=\{\mathbf{s}_{i}\}_{i=1}^{n}$ with node features $\mathbf{s}_{i}\in\mathbb{R}^{2}$, represented as a sequence of $n$ cities in a two-dimensional space. We are concerned with finding a permutation of the nodes $\pi=(\pi_{1},\pi_{2},\pi_{3},\dots,\pi_{n})$, termed a tour, that visits each city once and has the minimum total length. We define the length of a tour defined by a permutation $\pi$ as
$\mbox{minimize}\quad L(\pi|S)=\left\|\mathbf{s}_{\pi_{n}}-\mathbf{s}_{\pi_{1}}\right\|+\sum_{i=1}^{n-1}\left\|\mathbf{s}_{\pi_{i}}-\mathbf{s}_{\pi_{i+1}}\right\|,$ (15)